Assuming 10% as the annual rate of discount or interest, calculate the net present value for the following project:

Initial investment (an outflow) $ 40000

Revenue yr 1: $ 80000                                   Costs yr 1: $ 60000

Revenue yr 2: $100000                                   Costs yr2: $60000

Revenue yr 3: $120000                                   Costs yr 3: $ 70000
Answer #1

Let the cash Flows in Year n be denoted by CFn

Cash Flow = Revenue - Costs

CF0 = -40000

CF1 = 80000 - 60000 = 20000

CF2 = 100000 - 60000 = 40000

CF3 = 120000 - 70000 = 50000

Interest Rate =r = 10%

Hence, NPV = Sum of present value of future cash flows =  CF0 + CF1/(1+r) + CF2/(1+r)2 + CF3/(1+r)3

= -40000 + 20000/(1+0.10) + 40000/(1+0.10)2 + 50000/(1+0.10)3

= $48,805.41